Can someone help me with this problem?
-3/w+2=w/w+2

Found 2 solutions by haiderk, stanbon:
Answer by haiderk(1)

-3/w+2=w/w+2
-3(w+2)sq /(w+2)=w(w+2)sq /(w+2)
-3(w+2)=w(w+2)
-3w-6=wsq + 2w
wsq + 5w +6 = 0
wsq +2w + 3w + 6= 0
w(w+2) +3 (w+2) = 0
(w+3)(w+2)=0
W= -3 W=-2

Answer by stanbon(75887)

It looks like w must be -3 but it's not
possible to tell what the denominator is
on your posted problem. Please use parentheses
where appropriate.
Cheers,
Stan H.